OpenStudy (anonymous):

Find the product. (3v+4)(3v-4)

OpenStudy (mertsj):

Use this property: (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

oh duh lol, so 9v^2-16?

OpenStudy (mathstudent55):

Correct. If you forget that (a + b)(a - b) = a^2 - b^2, you can always use FOIL.

OpenStudy (anonymous):

ah yes! Thanks

OpenStudy (mathstudent55):

wlcm